# Minimax discrete distribution estimation with self-consumption

### v1

@@ Abstract
Learning distributions from i.i.d. samples is a well-understood problem. However, advances in generative machine learning prompt an interesting new, non-i.i.d. setting: after receiving a certain number of samples, an estimated distribution is fixed, and samples from this estimate are drawn and introduced into the sample corpus, undifferentiated from real samples. Subsequent generations of estimators now face contaminated environments, an effect referred to in the machine learning literature as self-consumption. In this paper, we study the effect of such contamination from previous estimates on the minimax loss of multi-stage discrete distribution estimation.
In the data accumulation setting, where all batches of samples are available for estimation, we provide minimax bounds for the expected $\ell_{2}^{2}$ and $\ell_{1}$ losses at every stage. We show examples where our bounds match under mild conditions, and there is a strict gap with the corresponding oracle-assisted minimax loss where real and synthetic samples are differentiated. We also provide a lower bound on the minimax loss in the data replacement setting, where only the latest batch of samples is available, and use it to find a lower bound for the worst-case loss for bounded estimate trajectories.
@@ I Introduction
A sketch of the data production workflow considered in our paper is shown in Figure 1. A crucial aspect of the model is that the corpus does not include the identities of the source of the samples, i.e. whether they are real or synthetic. In the data accumulation setting, new samples are continually added to the corpus at every stage, whereas it consists of only fresh samples in the data replacement setting. This mirrors quite a few variants considered in the generative machine learning literature. In this paper, we assume that each sample is drawn from either the true underlying distribution or the previous estimate based on the outcome of a coin flip with time-dependent bias $\alpha_{t}$, assumed to be known to the estimator.
Figure 1: The estimation workflow under self-consumption
I-A Related work
@@
Model collapse in settings containing feedback loops with different mechanisms of mixing synthetic data with old and new real samples was first studied empirically in [5] and [6]. Specifically, the case for discrete distribution estimation in a fully synthetic loop, wherein, at every stage, the estimator only has access to samples generated from the distribution estimated at the previous stage was considered as a motivating example in the analysis of [5].
@@
An upper bound on the proportion of synthetic data for in-loop stability of parametric generative models trained on freshly generated (mixed) samples at each stage was provided in [7]. The evolution of parametric models and nonparametric kernel density estimators trained on fully synthetic and mixed data loops was described in [8]. Model collapse in regression models and mitigation strategies are discussed in [9, 10].
I-B Contributions
In this work, we pose the multi-stage distribution estimation problem with self-consumption of samples, which has not been studied in the minimax setting before in the literature. Moreover, we provide the first (to the best of our knowledge) minimax lower bounds on the $\ell_{2}^{2}$ and $\ell_{1}$ losses at each stage in terms of properties of the estimates at previous stages in Theorem 1 (for the data accumulation workflow) and Corollary 3 (for the data replacement workflow).
We also propose a sequence of estimators that is order optimal under certain regimes with stage-wise performance guarantees in Theorem 2, and show examples of problem parameters where the conditions for order optimality are met in the data accumulation workflow. We also show that the ratio of minimax loss to the minimax loss in the presence of identity information grows infinitely large in such cases.
For example, when the number of samples added in stage $t$ increases linearly with $t,$ but each stage only contains, on average, $\sqrt{t}$ samples from the true distribution, then we show that the loss under self-consumption behaves as $1/\sqrt{t}$ while the oracle bound, with known sample identities, would behave as $1/t.$
Finally, we use our lower bound for the data replacement workflow to find a lower bound on the worst-case loss given a time-independent upper bound.
@@ I-C Notation
Deterministic quantities and functions are represented by lowercase Greek and Roman alphabets; random variables are represented by uppercase Roman alphabets. The probability simplex corresponding to the alphabet of size $k$ is denoted as $\Delta_{k}$. The $j^{\mathrm{th}}$ component of a discrete distribution $P$ or vector $p$ is indexed by square brackets as $P[j]$ and $p[j]$ respectively, and the probability of an event $A$ under distribution $Q$ is denoted as $Q\{A\}$. The first subscript of quantities in sequences denotes the index in the sequence, subsequent ones are used to denote other parameters when needed. The $n$-fold product of a distribution $Q$ is denoted as $Q^{n}$. For $t\geq 0$ and functions $f,g$, we write $g(t)=\Theta(f(t))$ if there exist constants $a_{1},a_{2}$ such that $a_{1}f(t)\leq g(t)\leq a_{2}f(t)\ \forall t\geq 0$. We use $[a:b]$ to denote $\{x\in\mathbb{N}:a\leq x\leq b\}$. The unit vector in with component $j$ equal to $1$ is denoted as $e_{j}$.
@@ II Problem Setup
Definition 1 (Self-consuming estimation)
For a distribution $p\in\Delta_{k}$, a self-consuming sequence of sample-estimator pairs with sample sizes $(n_{t})_{t\geq 0}$ and sampling probabilities $(\alpha_{t})_{t\geq 1}$ is a sequence $\left(X_{t}^{n_{t}},\hat{p}_{t}(\cdot)\right)_{t\geq 0}$ such that
@@
The joint distribution of the random vector $(X_{0}^{n_{0}},\dots,X_{t}^{n_{t}})$ is denoted as $\bar{\mathbf{P}}_{t,({p,\{n_{i},\alpha_{i}:i\geq 0\}})}$. We drop the sample sizes and sampling probabilities when they are obvious from the context. By convention, $\alpha_{0}=1$.
The definition of this sequence encapsulates the following process: at stage 00, an estimate $\hat{P}_{0}$ is produced using a batch of $n_{0}$ i.i.d. samples from $p$. At every subsequent stage $t$, a batch of $n_{t}$ samples is produced, with each sample being distributed as $p$ with probability $\alpha_{t}$ and as $\hat{P}_{t-1}$ with probability $1-\alpha_{t}$.
For a sequence of estimators $\hat{p}_{t}(\cdot)$, we consider the minimax loss under the $\ell_{2}^{2}$ and $\ell_{1}$ loss metrics.
Definition 2 (Minimax loss)
For a given loss metric $\ell$, sequences $\{n_{i},\alpha_{i}:i\geq 0\}$ and fixed estimators $\hat{p}_{i}(\cdot),i\in[1:t]$, the minimax loss at stage $t$ is defined as
@@
It is important to note that the minimax loss is defined for the estimator at every stage $t$ while fixing the estimators in all the previous stages. This follows from the process of designing the estimators: they are selected sequentially at every stage, without concern for how the samples generated from these estimates might influence future samples.
@@ III Main Results
The main contributions of this paper are upper and lower bounds on the minimax loss of the self-consuming sequence at an arbitrary stage $t\geq 0$. The bounds are stated as follows:
Theorem 1 (Minimax lower bound)
Fix $\hat{p}^{(i)},i{\in}[0{:}t{-}1]$. Let $g_{0}(\epsilon)=0\ \forall\epsilon$. For $i\neq 0$, let
$g_{i}(\epsilon):=\sup_{p}\max_{j}\bar{\mathbf{P}}_{i-1,({p})}\left\{\hat{P}_{i-1}[j]<p[j]-\epsilon\right\}$
@@
be the worst case probability of an $\epsilon$-error under $\bar{\mathbf{P}}_{i,({p})}$. Denote $h_{t}(\nicefrac{{1}}{{4k}}):=\sum_{i=0}^{t}n_{i}\alpha_{i}g_{i}(\nicefrac{{1}}{{4k}})$.
@@
The minimax loss for the self-consuming estimator-sample sequence at stage $t$ under $\ell_{2}^{2}$ and $\ell_{1}$ losses is lower bounded, respectively, as
$r^{\ell_{2}^{2}}_{t,k}\geq$ $\frac{\nicefrac{{1}}{{512}}}{\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}+h_{t}({\nicefrac{{1}}{{4k}}})}$ (1)
$r^{\ell_{1}}_{t,k}\geq$ $\sqrt{\frac{\nicefrac{{k}}{{4096}}}{\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}+h_{t}({\nicefrac{{1}}{{4k}}})}}$ (2)
whenever $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}+h_{1/4k}(t)\geq k/4$.
Theorem 2 (Upper bound)
@@
Let $n_{i}\alpha_{i}\leq\sum_{j=0}^{i}n_{j}\alpha_{j}^{2}$ for every $i\leq t$. Then there exists an estimator-sample sequence for which the $\ell_{2}^{2}$ and $\ell_{1}$ losses are upper bounded as
$\sup_{p}E[\ell_{2}^{2}(p,\hat{p}_{t}(X_{0}^{n_{0}},\dots,X_{t}^{n_{t}}))]\leq$ $\frac{1-\nicefrac{{1}}{{k}}}{\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}}$ (3)
$\sup_{p}E[\ell_{1}(p,\hat{p}_{t}(X_{0}^{n_{0}},\dots,X_{t}^{n_{t}}))]\leq$ $\sqrt{\frac{k-1}{\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}}}$ (4)
@@
Theorem 1 can also be modified for the data replacement setting to obtain the following corollary:
Corollary 3 (Minimax lower bound with data replacement)
@@
Let $\mathbf{Q_{p}}$ be a distribution on $\Delta_{k}$ that depends on $p$, and let $\hat{P}\sim\mathbf{Q_{p}}$ be an estimate of $p$. Let samples $X_{1}^{n_{1}}\sim\tilde{P}({p,\hat{P},\alpha})^{n_{1}}$. Denote $h({\epsilon}):=n_{1}\alpha\sup_{p}\max_{j}\mathbf{Q_{p}}\{\hat{P}[j]<p[j]-\epsilon\}$. The minimax $\ell_{2}^{2}$ and $\ell_{1}$ losses of an estimator $\hat{p}_{1}(X_{1}^{n_{1}})$ are lower bounded as
@@
whenever $n_{1}\alpha^{2}+h\left({\nicefrac{{1}}{{4k}}}\right)\geq k/4$.

### v2

# Model non-collapse: Minimax bounds for recursive discrete distribution estimation

@@ Abstract
Learning discrete distributions from i.i.d. samples is a well-understood problem. However, advances in generative machine learning prompt an interesting new, non-i.i.d. setting: after receiving a certain number of samples, an estimated distribution is fixed, and samples from this estimate are drawn and introduced into the sample corpus, undifferentiated from real samples. Subsequent generations of estimators now face contaminated environments, a scenario referred to in the machine learning literature as self-consumption.
Empirically, it has been observed that models in fully synthetic self-consuming loops collapse—their performance deteriorates with each batch of training—but accumulating data has been shown to prevent complete degeneration. This, in turn, begs the question: What happens when fresh real samples are added at every stage? In this paper, we study the minimax loss of self-consuming discrete distribution estimation in such loops. We show that even when model collapse is consciously averted, the ratios between the minimax losses with and without source information can grow unbounded as the batch size increases.
In the data accumulation setting, where all batches of samples are available for estimation, we provide minimax lower bounds and upper bounds that are order-optimal under mild conditions for the expected $\ell_{2}^{2}$ and $\ell_{1}$ losses at every stage. We provide conditions for regimes where there is a strict gap in the convergence rates compared to the corresponding oracle-assisted minimax loss where real and synthetic samples are differentiated, and provide examples where this gap is easily observed. We also provide a lower bound on the minimax loss in the data replacement setting, where only the latest batch of samples is available, and use it to find a lower bound for the worst-case loss for bounded estimate trajectories.
@@ I Introduction
A sketch of the data production workflow considered in our paper is shown in Figure 1. A crucial aspect of the model is that the corpus does not include the identities of the source of the samples, i.e. whether they are real or synthetic. In the data accumulation setting, new samples are continually added to the corpus at every stage, whereas it consists of only fresh samples in the data replacement setting. This mirrors quite a few variants considered in the generative machine learning literature. In this paper, we assume that each sample is drawn from either the true underlying distribution or the previous estimate based on the outcome of a coin flip, independent of every other sample, with stage-dependent bias $\alpha_{t}$ which is assumed to be known to the estimator.
Figure 1: The estimation workflow in the data accumulation setting; solid arrows indicate sampling (with the corresponding sampling probability indicated next to them), whereas dotted arrows indicate estimation. Under the data replacement setting, the corpus used for estimation only contains the latest batch of samples.
I-A Contributions
In this work, we pose the multi-stage distribution estimation problem with self-consumption of samples, which has not been studied in the minimax setting before in the literature. We analyze the minimax loss in terms of the sequences $n_{t}$, denoting the total number of samples in batch $t$, and $\alpha_{t}$, the probability with which each sample in batch $t$ is drawn from the real distribution. We consider the oracle-assisted minimax loss—where the source distribution of each sample is provided to the estimator by an oracle—as a baseline. It is perhaps unsurprising that the oracle-assisted minimax loss at stage $t$ is $\Theta(\nicefrac{{1}}{{\sum_{i=0}^{t}n_{i}\alpha_{i}}})$, since the expected number of real samples seen after $t$ batches is precisely $\sum_{i=0}^{t}n_{i}\alpha_{i}$.
We provide the first (to the best of our knowledge) information-theoretic lower bounds on the $\ell_{2}^{2}$ and $\ell_{1}$ losses at each stage in terms of properties of the estimates at previous stages in Theorem 1 (for the data accumulation workflow) and Corollary 4 (for the data replacement workflow). We also propose a sequence of estimators that is order-optimal under certain regimes with stage-wise performance guarantees in Theorem 2. We show conditions where there are gaps between the upper bound and lower bounds on the minimax loss and the oracle-assisted minimax loss the data accumulation workflow (Propositions 2 and 3) and provide examples where these conditions are satisfied. Finally, we use our lower bound for the data replacement workflow to find a lower bound on the worst-case loss given a time-independent upper bound.
We find that the minimax loss stays within a constant factor of the oracle-assisted one when $\alpha_{t}$ is lower bounded by a constant. When $\alpha_{t}$ decays to 00, i.e. when synthetic sampling outpaces real sampling, the effective sample size of batch $t$ is proportional to $n_{t}\alpha_{t}^{2}$ (as opposed to simply $n_{t}\alpha_{t}$ for the oracle-assisted setting) in most practically conceivable conditions. As a consequence, the ratio of the minimax loss to the oracle-assisted minimax loss can grow unbounded with $t$ (e.g. Claim 1). We specify the conditions for this regime in Proposition 2; the basic requirement is that $\alpha_{t}$ does not decay too quickly.
When $\alpha_{t}$ does decay quickly, an error term starts dominating the lower bound, and there is a gap between the upper and lower bounds. However, under mild conditions, there is also a gap between the lower bound and oracle-assisted loss (Claim 2): the effective sample size for the lower bound becomes proportional to $\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}}$; it remains $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$ for the upper bound. This regime is analyzed in detail in Proposition 3.
The various regimes are summarized in Table I.
TABLE I: The various regimes of non-collapsing discrete distribution estimation under $\ell_{2}^{2}$ loss in the accumulation workflow
Regime Upper Bound (Theorem 2) Lower Bound (Theorem 1) Oracle-assisted loss (Lemma 5)
Conditions on $\alpha_{t}$ Other conditions
$\alpha_{i}=\Omega(1)$ (Proposition 1) Both minimax losses are $\Theta\left(\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}}\right)$
$\alpha_{t}\downarrow 0$, but $\alpha_{t}=\Omega\left(\frac{1}{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}\right)$ (Proposition 2) Bounds match, minimax loss is $\Theta\left(\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}}\right)$ $\Theta\left(\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}}\right)$
$\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}}=o\left(\alpha_{t}\right)$ Additionally, the ratio of the upper and lower bounds is bounded by a constant independent of $k$ as $t\uparrow\infty$
$\alpha_{t}\downarrow 0$ and $\alpha_{t}=o\left(\frac{1}{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}\right)$ $n_{t}\alpha_{t}\leq\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$, $n_{t}\alpha_{t}=\mathcal{O}\left(\sum_{i=0}^{t-1}n_{i}\alpha_{i}\right)$ and $\sum_{i=0}^{t}n_{i}\alpha_{i}\uparrow\infty$ $\mathcal{O}\left(\frac{1}{\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}}\right)$ (Proposition 3) $\Omega\left(\frac{1}{\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}}}\right)$
$n_{t}\alpha_{t}=\mathcal{O}\left({\sum_{i=0}^{t-1}n_{i}\alpha_{i}}\right)$ and $\sum_{i=0}^{t}n_{i}\alpha_{i}\uparrow\infty$ (Corollary 3) $\mathcal{O}\left(1\right)$
${\sum_{i=0}^{t-1}n_{i}\alpha_{i}}=o(n_{t}\alpha_{t})$ $\Omega\left(\frac{1}{n_{t}\alpha_{t}}\right)$ $\Theta\left(\frac{1}{n_{t}\alpha_{t}}\right)$
(Special case) $\sum_{i=0}^{t}n_{i}\alpha_{i}=\mathcal{O}(1)$ All minimax losses are $\Theta\left(1\right)$
A few additional technical conditions are required for the results shown in the table: The lower bound on the loss at stage $t$ in Proposition 3 assumes the existence of estimators matching the lower bounds up to a universal constant factor for stages 00 through $t-1$ and the lower bounds in general only hold when $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}+h_{t}(\nicefrac{{1}}{{4k}})\geq k/4$. An elementary upper bound when the hypothesis conditions for Theorem 2 are not met is given in Corollary 3.
I-B Related work
@@
Model collapse in settings containing feedback loops with different mechanisms of mixing synthetic data with old and new real samples was first studied empirically in [5] and [6]. Specifically, the case for discrete distribution estimation in a fully synthetic loop, wherein, at every stage, the estimator only has access to samples generated from the distribution estimated at the previous stage was considered as a motivating example in the analysis of [5]. Further progress centered around collapse in fully synthetic loops in discrete distribution estimation was made in [7, 8].
@@
There has been a significant amount of recent work on model collapse in synthetic and mixed data loops. An upper bound on the proportion of synthetic data for in-loop stability of parametric generative models trained on freshly generated (mixed) samples at each stage was provided in [9]. The evolution of parametric models and nonparametric kernel density estimators trained on fully synthetic and mixed data loops was described in [10]. Model collapse in regression models and model collapse mitigation strategies are discussed in [11, 12, 13]; a lower bound on the expected error for regression models that do not account for self-consumption was presented in [14]. A stochastic-approximation-style treatment showing eventual degeneration for accumulating loops with no fresh real data is provided in [15]. Continuing progress on averting model collapse in data accumulation-style workflows, upper bounds on the estimation error in self-consuming loops where real data remains in the training set are studied in [16, 17, 18, 19, 20].
@@ I-C Notation
Deterministic quantities and functions are represented by lowercase Greek and Roman alphabets; random variables are represented by uppercase Roman alphabets. The probability simplex corresponding to the alphabet of size $k$ is denoted as $\Delta_{k}$. We use $[a:b]$ to denote $\{x\in\mathbb{N}:a\leq x\leq b\}$. The unit vector in with component $j$ equal to $1$ is denoted as $e_{j}$.
The $j^{\mathrm{th}}$ component of a discrete distribution $P$ or vector $p$ is indexed by square brackets as $P[j]$ and $p[j]$ respectively, and the probability of an event $A$ under distribution $Q$ is denoted as $Q\{A\}$. The first subscript of quantities in sequences denotes the index in the sequence, subsequent ones are used to denote other parameters when needed. The $n$-fold product of a distribution $Q$ is denoted as $Q^{n}$.
For $t\geq 0$ and functions $f,g$, we write $g(t)=\Theta(f(t))$ if there exist constants $a_{1},a_{2}\geq 0$ such that $a_{1}f(t)\leq g(t)\leq a_{2}f(t)\ \forall t\geq 0$; $g(t)=\mathcal{O}(f(t))$ if there exists a constant $a_{3}\geq 0$ such that $g(t)\leq a_{3}f(t)\ \forall t\geq 0$; $g(t)=\Omega(f(t))$ if there exists a constant $a_{4}>0$ such that $g(t)\geq a_{4}f(t)\forall t\geq 0$; $g(t)=o(f(t))$ if $\lim_{t\uparrow\infty}g(t)/f(t)=0$.
@@ II Problem Setup
We consider the problem of recursive discrete distribution estimation over finite alphabets. Let $\mathcal{X}:=\{1,2,\dots,k\}$ be the alphabet of size $k$. A sequence of sample-estimator pairs is defined as follows.
Definition 1 (Self-consuming estimation).
For a distribution $p\in\Delta_{k}$, a self-consuming sequence of sample-estimator pairs with sample sizes $(n_{t})_{t\geq 0}$ and sampling probabilities $(\alpha_{t})_{t\geq 1}$ is a sequence $\left(X_{t}^{n_{t}},\hat{p}_{t}(\cdot)\right)_{t\geq 0}$ of samples $X_{i}^{n_{i}}\in\mathcal{X}^{n_{i}}$ and estimators $\hat{p}_{i}:\mathcal{X}^{n_{0}}\times\dots\times\mathcal{X}^{n_{t}}\rightarrow\Delta_{k}$ such that
@@
The joint distribution of the random vector $(X_{0}^{n_{0}},\dots,X_{t}^{n_{t}})$ is denoted as $\bar{\mathbf{P}}_{t,({p,\{n_{i},\alpha_{i}:i\geq 0\}})}$. We drop the sample sizes and sampling probabilities from the subscript when they are obvious from the context. By convention, $\alpha_{0}=1$.
The definition of this sequence encapsulates the following process: at stage 00, an estimate $\hat{P}_{0}$ is produced using a batch of $n_{0}$ i.i.d. samples from $p$. At every subsequent stage $t+1$, a batch of $n_{t+1}$ samples is produced, with each sample being distributed as $p$ with probability $\alpha_{t+1}$ and as $\hat{P}_{t}$ with probability $1-\alpha_{t+1}$. Since this process can be seen as conditional sampling based on the outcome of a biased coin-flip, we refer to the outcome of this flip as the ‘identity’ of the source.
For a sequence of estimators $\hat{p}_{t}(\cdot)$, we consider the minimax loss under the $\ell_{2}^{2}$ and $\ell_{1}$ losses.
Definition 2 (Minimax loss).
For a given loss function $\ell:\Delta_{k}\times\Delta_{k}\rightarrow\mathbb{R}_{+}$, sequences $\{n_{i},\alpha_{i}:i\geq 0\}$ and fixed estimators $\hat{p}_{i}(\cdot),i\in[1:t]$, the minimax loss at stage $t$ is defined as
@@
It is important to note that the minimax loss is defined for the estimator at every stage $t$ while fixing the estimators in previous stages. This follows from the process of designing the estimators: they are selected sequentially, optimized at every stage, without concern for how the samples generated from these estimates might influence future samples and without mechanisms for modifying previously made estimates.
@@ III Main Results
The main contributions of this paper are upper and lower bounds on the minimax loss of the self-consuming sequence at an arbitrary stage $t\geq 0$. Recall that at stage $t$, the probability of drawing a sample from the true distribution $p$ is denoted by $\alpha_{t}$, the estimated distribution is denoted by $\hat{P}_{t}$, and the joint distribution of the samples collected and distributions estimated so far is denoted by $\bar{\mathbf{P}}_{t,(p)}$.
III-A Bounds on the minimax loss
The lower bounds on the minimax loss at stage $t$ are given in terms of the performance of the previous estimates, encapsulated in the error term $h_{t}(\epsilon)$, defined as follows:
Definition 3.
The error term $h_{t}(\epsilon)$ is defined as
$h_{t}(\epsilon):=\sum_{i=0}^{t}n_{i}\alpha_{i}g_{i}(\epsilon),$
where
$g_{i}(\epsilon)$ $:=\sup_{p}\max_{j}\bar{\mathbf{P}}_{i-1,({p})}\left\{\hat{P}_{i-1}[j]<p[j]-\epsilon\right\}$ $i\neq 0$
@@ III-A Bounds on the minimax loss
denotes the worst case (over component and distribution) probability of an $\epsilon$-error in $\hat{P}_{i-1}$ under $\bar{\mathbf{P}}_{i-1,({p})}$, and $g_{0}(\epsilon)=0\ \forall\epsilon$.
We are now ready to state the main results of this paper.
Theorem 1 (Minimax lower bound).
@@
Fix $\hat{p}^{(i)},i{\in}[0{:}t{-}1]$. The minimax loss for the self-consuming estimator-sample sequence at stage $t$ under $\ell_{2}^{2}$ and $\ell_{1}$ losses is lower bounded, respectively, as
$r^{\ell_{2}^{2}}_{t,k}\geq$ $\frac{\nicefrac{{1}}{{512}}}{\sum\limits_{i\in[0:t]}n_{i}\alpha_{i}^{2}\,+h_{t}({\nicefrac{{1}}{{4k}}})}$ (1)
$r^{\ell_{1}}_{t,k}\geq$ $\sqrt{\frac{\nicefrac{{k}}{{4096}}}{\sum\limits_{i\in[0:t]}n_{i}\alpha_{i}^{2}\,+h_{t}({\nicefrac{{1}}{{4k}}})}},$ (2)
whenever $\sum\limits_{i\in[0:t]}n_{i}\alpha_{i}^{2}\,+h_{t}(\epsilon)\geq k/4$.
Theorem 2 (Upper bound).
@@
Let $n_{i}\alpha_{i}\leq\sum_{j\in[0:i]}n_{j}\alpha_{j}^{2}$ for every $i\leq t$. Then there exists a sequence of estimators $\hat{p}_{i},i\geq 0$ for which the worst-case $\ell_{2}^{2}$ and $\ell_{1}$ losses are upper bounded as
$\sup_{p}E[\ell_{2}^{2}(p,\hat{p}_{t}(X_{0}^{n_{0}},\dots,X_{t}^{n_{t}}))]\leq$ $\frac{1-\nicefrac{{1}}{{k}}}{\sum\limits_{i\in[0:t]}n_{i}\alpha_{i}^{2}}$ (3)
$\sup_{p}E[\ell_{1}(p,\hat{p}_{t}(X_{0}^{n_{0}},\dots,X_{t}^{n_{t}}))]\leq$ $\sqrt{\frac{k-1}{\sum\limits_{i\in[0:t]}n_{i}\alpha_{i}^{2}}}.$ (4)
We can directly observe that when the error term $h_{t}(\nicefrac{{1}}{{4k}})$ is small, the minimax errors are within constant factors of each other; and, as we will see in the baseline, the erasure of the source information leads to a penalty of $\alpha_{t}$ on the effective sample size of batch $t$ (as compared to when the estimators know which distribution each sample is drawn from). As a corollary of the analysis of Theorem 2, we also describe the worst-case error when the sequence of empirical estimators without further processing is used. This can be used in place of Theorem 2 if the hypothesis conditions are not satisfied, and might also be of independent interest.
Corollary 3 (Unprocessed upper bound).
For the sequence of estimators $\hat{P}_{t}^{\mathrm{emp}}:=\hat{p}_{\mathrm{emp}}(X_{0}^{n_{0}},\dots,X_{t}^{n_{t}})$, where $\hat{p}_{\mathrm{emp}}(\cdot)$ denotes the empirical estimator of all samples, the worst-case $\ell_{2}^{2}$ loss under self-consumption at stage $t$ is bounded above as
$\sup_{p\in\Delta_{k}}E[\ell_{2}^{2}(\hat{P}_{t}^{\mathrm{emp}},p)]\leq\left(1-\frac{1}{k}\right)\cdot\sum_{i=0}^{t}\frac{n_{i}}{\left(\sum_{j=0}^{i}n_{j}\right)^{2}}$ (5)
We note that this bound increases with $t$, and therefore it is perhaps useful only to use the upper bound $\nicefrac{{1}}{{n_{0}}}$ obtained by using the order-optimal estimator for just the $0^{\mathrm{th}}$ batch. However, it is interesting to note that in both cases, there exists a constant upper bound independent of $t$—the sum in the RHS of (5) is exactly $\nicefrac{{1}}{{n_{0}}}$ plus a (lower) Riemann sum of the function $\nicefrac{{1}}{{x^{2}}}$ from $n_{0}$ to $\sum_{i=0}^{t}n_{i}$ with intervals $n_{i}$. Therefore, the upper bound on the minimax error stays $\mathcal{O}(1)$. In fact, if each batch is the same size $n$, we find an upper bound of $\nicefrac{{\pi^{2}}}{{6}}\cdot\nicefrac{{1}}{{n}}$, recovering the additional penalty factor of $\nicefrac{{\pi^{2}}}{{6}}$ highlighted by [12, 16].
@@
In addition, for application in the data replacement workflow, Theorem 1 can be modified to obtain the following corollary:
Corollary 4 (Minimax lower bound with data replacement).
@@
Let $\mathbf{Q}_{p}$ be a distribution on $\Delta_{k}$ that depends on $p$, and let $\hat{P}\sim\mathbf{Q}_{p}$ be an estimate of $p$. Let samples $X_{1}^{n_{1}}\sim\tilde{P}({p,\hat{P},\alpha})^{n_{1}}$. The minimax $\ell_{2}^{2}$ and $\ell_{1}$ losses of an estimator $\hat{p}_{1}(X_{1}^{n_{1}})$ are lower bounded as
@@
whenever $n_{1}\alpha^{2}+h\left({\nicefrac{{1}}{{4k}}}\right)\geq k/4$; $h({\epsilon})$ here is defined as $n_{1}\alpha\sup_{p}\max_{j}\mathbf{Q}_{p}\{\hat{P}[j]<p[j]-\epsilon\}$.
The proofs of Theorem 1 and Corollary 4 are deferred to Appendix A and the proofs of Theorem 2 and Corollary 3 are deferred to Appendix B. Since the bounds for $\ell_{1}$ loss are proportional to $\sqrt{k}$ times the square root of the $l_{2}^{2}$ bounds, we only consider $\ell_{2}^{2}$ bounds in the rest of the paper.
III-B Baseline: oracle-assisted minimax loss
To understand the impact of not knowing the source identity on the minimax loss, it is useful to compare the upper and lower bounds presented above to the loss in the case where the identity of the source of each sample is known in the form of an auxiliary random vector $W:=W_{i}^{n_{i}}\sim(\mathrm{Bern}(\alpha_{i}))^{n_{i}},i\in\{1,\dots,t\}$ provided to the estimator by an oracle such that, conditioned on whether $W_{i}[s]$ is 00 or $1$, the sample $X_{i}[s]$ is distributed either as $p$ or $\hat{P}_{i-1}$ respectively.
Lemma 5 (Oracle-assisted minimax loss).
The minimax loss $r^{\ell_{2}^{2},\mathrm{oracle}}_{t,k}$ of self-consuming estimation with information on source identity is bounded between
$\frac{\nicefrac{{1}}{{512}}}{\sum\limits_{i\in[0:t]}n_{i}\alpha_{i}}\leq r^{\ell_{2}^{2},\mathrm{oracle}}_{t,k}\leq\frac{2}{\sum\limits_{i\in[0:t]}n_{i}\alpha_{i}}.$
The proof of this Lemma is deferred to Appendix C.